Will you help me with substitution 3x+y=24 y=3x
Answer by Iluvidina(15)

Y=12; x=4
I started with y=3x, because it is the easiest to figure out what one of the variables equals
I divided each side by 3. The equation then was 1/3y=x
Then, I substituted that in the other equation for x
The equation 3x+y=24 became 3(1/3y)+y=24
Then you multiply 3 by 1/3y, becoming (1)y+y=24
y+y becomes 2y, so the equatiion becomes 2y=24
Then you divide each side by 2, becoming y=12
Then you substitute y for 12 in the equation y=3x. So, it would be 12=3x.
Then you divide 3 on both sides, becoming x=4